OpenStudy (anonymous):

*help requested* *fan and medal awarded* Vivian combined different amounts of white paint, blue paint, and green paint to make 145 milliliters (ml) of paint of a new shade. The amount of white paint was x ml. The amount of blue paint was 10 ml less than twice the amount of white paint, and the amount of green paint was half the amount of blue paint. This is represented in the equation below. x + (2x – 10) + 1/2 (2x – 10) = 145 What was the difference in the amounts of white paint and blue paint Vivian combined? 30 ml 35 ml 40 ml 45 ml

OpenStudy (anonymous):

Just solve the eqn then you have 'x' i.e. white paint and can simply workout how much blue paint.

OpenStudy (anonymous):

uhmmmmm

OpenStudy (anonymous):

okay i think its b. am i right ?

OpenStudy (anonymous):

SHould get: x+2x+x-10-5=145

OpenStudy (anonymous):

Now if you can get the like term together you can work out the solution.

OpenStudy (anonymous):

the like term ? okay im not good at math at all so can someone either explain it very simply or jus explain the whole with the answer ????????????

OpenStudy (anonymous):

get all the 'x' s together and get the -10 and -5 together. If you're having problems with this you have gaps in your knowledge.

OpenStudy (anonymous):

dont insult me jus help me or i will report u

OpenStudy (anonymous):

x+2x+x = ? -10-5 = ?

OpenStudy (anonymous):

have to go, try yourself will be back in a few hours. ;D

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

Are you in connections

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh I could have helped because I did an assignment with those questions.

OpenStudy (anonymous):

thanks can u explain ?

OpenStudy (anonymous):

No I dont even know how I got the right answer. I was just guessing.

OpenStudy (anonymous):

well what did u guess ?

OpenStudy (anonymous):

@Piggiepigpigs